OpenStudy (anonymous):

x2 + y2 = 13 2x - y = 4

OpenStudy (sburchette):

System of equations I assume?

OpenStudy (anonymous):

yes UGH!! quadratic systems...I hate this stuff

OpenStudy (anonymous):

1) x^2 + y^2 = 13 2) 2x - y =4 rearrange the second one to be y=ax + b y = 2x + 4 sub y = 2x + 4 into the first equation x^2 + (2x + 4)^2 = 13 x^2 + 4x^2 + 16x + 16 = 13 5x^2 + 16x + 16 - 13 =0 5x^2 + 16x + 3 = 0 (5x + 1)(x+3) if u have to solve further. 5x + 1 = 0 x + 3 = 0 5x = -1 x = -3 x = -1/5

OpenStudy (sburchette):

All right, it is easiest to solve for y in the second equation. This gives us\[y=2x-4\]You can substitute for y in the first equation and you get\[x^2+(2x-4)^2=13\]Expanding the (2x-4)^2 and collecting like terms gives us\[5x^2-16x+3=0\]We can use the quadratic formula with a=5 b=-16 and c=3\[\frac{16 \pm \sqrt{(-16)^2-4(5)(3)}}{2(5)} \]Simplifying we get, \[\frac{16 \pm 14}{10}\]Which is 3 or 1/5

OpenStudy (anonymous):

oh okay I was almost there I just was lost at the final equation

OpenStudy (anonymous):

follow @SBurchette, i mistakenly forgot the minus. sorry!

OpenStudy (anonymous):

the negative answers I got should be positive.

OpenStudy (anonymous):

okie dokie THANKS THE BOTH OF YOU!!!!!!!!

OpenStudy (sburchette):

No problem =)

OpenStudy (anonymous):

:)